# Simple Semantic-Aided Few-Shot Learning

### 20 Mar 2024 v2

@@ 1 Introduction
In this paper, we propose a simple framework called Semantic aided Few-shot learning (SemFew). To begin with, we identify the vital role of high-quality semantics in FSL and innovatively propose an automatic step-by-step Semantic Evolution (SemEvo) process to acquire detailed and accurate semantics. It first converts the class name into a short description that matches the image content of each class. Next, to further enrich recognizable characteristics, we expand and paraphrase short descriptions. The paraphrased descriptions contain more class-related knowledge than class names and short descriptions, making them better encompass a vast array of visual features. To illustrate, Fig. 1 showcases an instance that the paraphrased semantic contains a large number of detailed descriptions matching visual features e.g. organ characteristics and eating habitat. With this information, the model can better reconstruct an image with periphery feature (left image), into an image with prototype feature (right image), which benefits few-shot classification. Then, to fully exploit the advantage of paraphrased semantics, we design a network called Semantic Alignment Network (SemAlign). Different from previous methods, SemAlign does not adopt complex semantic understanding modules but utilizes a basic two-layer network. After accepting semantics and visual features as input, it reconstructs prototype features for subsequent classification. Experimental results on six widely adopted benchmarks demonstrate the effectiveness of our method, providing a simple insight into future FSL research. Our contributions can be summarized as three points:
@@
    To the best of our knowledge, we are the first to consider the automatic way of collecting high-quality semantics and applying them in few-shot learning.
@@ 2 Related Work
The significant issue in the field of Few-Shot Learning is how to effectively extract valuable information from limited data. Targeting this issue, existing research explored two different types of approaches, i.e., visual-based methods and semantic-based methods.
@@ 3.1 Background
where $d(\cdot,\cdot)$ means distance metric, e.g. cosine distance, and $N$ denotes the number of classes.
@@ 3.3 Semantic Alignment Network
where $r_{t}$ is the reconstructed prototype. To ensure the reconstructed prototype is further aligned with the ground-truth class prototype, it will be fused with support images as follows:
@@
where $p_{t}$ means the classification prototype of class $t$, $u_{t}=\frac{1}{K}\sum_{i=1}^{K}f(x_{i})$ is the mean vector of support samples, $r_{t}$ is the reconstructed prototype, and $k\in[0,1]$ is a manually controlled fusion factor. Then, Eq. 2 is utilized to assign the label of the query sample, where $d(\cdot,\cdot)$ is set as cosine distance.
@@ 5 Conclusion
To unleash the potential of semantics in few-shot learning, we propose the Semantic Evolution to automatically generate high-quality semantics from simple class names. Then, we design a simple Semantic Alignment Network to transform the concatenated visual features and high-quality semantics features into robust class prototypes for classification. The experimental results show that with high-quality semantics, the basic network can easily achieve greater performance compared to other state-of-the-art methods.

### 09 Apr 2024 v3

@@ 1 Introduction
In this paper, we propose a simple framework called Semantic-Aided Few-Shot Learning (SemFew). To begin with, we identify the vital role of high-quality semantics in FSL and innovatively propose an automatic step-by-step Semantic Evolution (SemEvo) process to acquire detailed and accurate semantics. It first converts the class name into a short description that matches the image content of each class. Next, to further enrich recognizable characteristics, we expand and paraphrase short descriptions. The paraphrased descriptions contain more class-related knowledge than class names and short descriptions, making them better encompass a vast array of visual features. To illustrate, Fig. 1 showcases an instance that the paraphrased semantic contains a large number of detailed descriptions matching visual features e.g. organ characteristics and eating habitat. With this information, the model can better reconstruct an image with periphery feature (left image), into an image with prototype feature (right image), which benefits few-shot classification. Then, to fully exploit the advantage of paraphrased semantics, we design a network called Semantic Alignment Network (SemAlign). Different from previous methods, SemAlign does not adopt complex semantic understanding modules but utilizes a basic two-layer network. After accepting semantics and visual features as input, it reconstructs prototype features for subsequent classification. Experimental results on six widely adopted benchmarks demonstrate the effectiveness of our method, providing a simple insight into future FSL research. Our contributions can be summarized as three points:
@@
    To the best of our knowledge, we are the first to consider the automatic way of collecting high-quality semantics and applying them in FSL.
@@ 2 Related Work
The significant issue in FSL is how to effectively and accurately extract class-related features from limited data. Targeting this issue, existing research explored two different types of approaches, i.e., visual-based methods and semantic-based methods.
@@ 3.1 Background
where $d(\cdot,\cdot)$ means distance metric, e.g., cosine distance, and $N$ denotes the number of classes.
@@ 3.3 Semantic Alignment Network
where $r_{t}$ is the reconstructed prototype. To ensure the reconstructed prototype is further aligned with the ground-truth class prototype, it will be fused with support images in a convex combination manner [57]:
@@
where $p_{t}$ means the classification prototype of class $t$, $u_{t}=\frac{1}{K}\sum_{i=1}^{K}f(x_{i})$ is the mean vector of support samples, $r_{t}$ is the reconstructed prototype produced by SemAlign, and $k\in[0,1]$ is a manually controlled fusion factor. Then, Eq. 2 is utilized to assign label for each query sample, where $d(\cdot,\cdot)$ is set as cosine distance.
@@ 5 Conclusion
To unleash the potential of semantics in FSL, we propose the Semantic Evolution to automatically generate high-quality semantics from simple class names. Then, we design a simple Semantic Alignment Network to transform the concatenated visual features and high-quality semantics features into robust class prototypes for classification. The experimental results show that with high-quality semantics, the basic network can easily achieve greater performance compared to other state-of-the-art methods.